A chemist prepares a sample of hydrogen bromide and finds that it occupies 216 ml at 64 c and 535 torr. what volume would it occupy at0 c at the same pressure? answer in units of ml


1
Expert's answer
2022-05-13T13:33:03-0400

PV=nRT

P1=P2=535 torr=71327.5 Pa

V1=216ml=0.000216 m3

T1=64 °C=337 K

T2=0°C=273 K

R=8.31 J/(mole K)

"n=\frac{P_1V_1}{RT_1}=\frac{71327.5x0.000216}{8.31x337}=0.0055 mole"

"V_2=\frac{nRT_2}{P_2}=\frac{0.0055x8.31x273}{71327.5}=0.000175 m^3=175 ml"